2+2cos2xcos5x=(sin2x)sin2x please slove this trigonomertry eruation 

Grade:11

1 Answers

Ayanava Datta
6 years ago
2+2cos2xcos5x=1-cos2x
sin25x+cos2x +2cos2xcos5x+cos2x=0
sin25x+(cos5x+cos2x)2=0..............(1)
x is real then equ1 is satisfy by those value os x for which
sin5x=0.............(2)
&
cos5x+cos2x=0...................(3)
from 2 5x=nR -x=nr\5.............(4)
from 3  cos(7x\2)=0 ….(p)&cos(3x\2)=0.........(q)
from  p    x=(2n+1)R\7................(5)
from q     x=(2n+1)r\3......................(6)
where n is an integer 
common solution to  4&5/6 are x=[+-]R,[+-]3R,[+-]5R,...............
X=(2m+1)R where m is an intrger
[R=pie]